OpenStudy (anonymous):

Write an exponential function to model this situation: a population of 420 animals decreases at an annual rate of 21%. Then predict the value of the function after 5 years (to the nearest whole number).

OpenStudy (anonymous):

P(t)=420×(.79)t P(5)=420×(.79)5=129

OpenStudy (anonymous):

@DragonFlame : Thank you~

OpenStudy (anonymous):

No problem :)